Find the wavelength of the spectral line that corresponds to a transition in hydrogen from the n = 6 state to the n = 3 state. In what part of the spectrum is this?

1
Expert's answer
2021-12-02T10:07:15-0500

"1/\lambda=R(1/n-1/m)=10973731.534\cdot(1/3-1/6)=1828955.26\ (m^{-1})\to"


"\lambda=1/1828955.26=5.47\cdot10^{-7}\ (m)=547\ (nm)" , visible radiation. Answer